OpenStudy (anonymous):

find a counter example to the conditional or to its converse(one is true and the is not). If x is divisible by 3, then x^2 is divisible by 3. The converse is if x^2 is divisible by 3, then x is divisible by 3.

OpenStudy (dumbcow):

there is no counterexample to conditional if x is divisible by 3 then say x = m/3, m is any integer then x^2 = m^2/3^2 = (m^2/3)/3 thus divisible by 3 For converse there are counterexamples: say x^2 = 12 then x = 2sqrt(3) -> not divisible by 3

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

your welcome